OpenStudy (anonymous):

How do you find the square roots of the fraction -9/16

OpenStudy (kc_kennylau):

Real roots or complex number?

OpenStudy (anonymous):

sqrt(-9/16)? You're going to end up with a complex number. sqrt(-9/16) = sqrt(9/16) * sqrt(-1) sqrt(9/16) = sqrt(9)/sqrt(16) = 3/4 therefore sqrt(-9/16) = (3/4)i

OpenStudy (kc_kennylau):

Or -(3/4)i because there are two roots

OpenStudy (anonymous):

Ah yes, I assumed it was a positive root.

OpenStudy (kc_kennylau):

@shneagle get it?

OpenStudy (anonymous):

That makes sense, thanks so much!

OpenStudy (anonymous):

How do you find real roots?

OpenStudy (kc_kennylau):

If the question is finding real roots then none exists

OpenStudy (anonymous):

Thanks!